# A randomized lattice rule without component-by-component construction

### v1

@@ 3.2. Error bounds
$\leq\|f\|_{\alpha,\boldsymbol{\gamma}}\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}(2\zeta(\alpha/\lambda))^{|u|}\right)^{\lambda-\delta},$
@@
so that, by applying Cauchy–Schwartz inequality, we obtain
@@
$\leq(1-\eta)^{r}\|f\|_{\alpha,\boldsymbol{\gamma}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{2}(2\zeta(2\alpha))^{|u|}\right)^{1/2}.$
@@
$e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ $\leq\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}(2\zeta(\alpha/\lambda))^{|u|}\right)^{\lambda-\delta}$
$\qquad+(1-\eta)^{r}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{2}(2\zeta(2\alpha))^{|u|}\right)^{1/2},$
@@
Let $M\in\mathbb{N}$, with $M\geq 2$, be given and assume that $M$ and $\eta\in(0,1)$ satisfy (3) for all $1/2\leq\lambda<\alpha$. Let
@@
Then $e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ decays with the order $M^{-\lambda+\delta-1/2}\log M$ for any $1/2<\lambda<\alpha$ and $0<\delta<\min(\lambda-1/2,1)$, which is arbitrarily close to $M^{-\alpha-1/2}$ when $\lambda\to\alpha-0$ and $\delta\to 0+0$. Moreover, $e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ is bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$.
@@
Let $\alpha>1/2$ be a real number and $\boldsymbol{\gamma}\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ be a sequence of non-negative weights, with each $\gamma_{j}$ bounded above by $1$. Furthermore, let $r,M\in\mathbb{N}$, with $M\geq 2$, be given. Assume that $M$ and $\eta\in(0,1)$ satisfy (3) for all $1/2\leq\lambda<\alpha$. Then it holds that
@@
$\leq\frac{rC_{\lambda,\delta}}{\alpha(1-\eta)^{\lambda(2-\delta)}M^{\lambda(2-\delta)+1}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}\left(2\zeta(\alpha/\lambda)\right)^{|u|}\right)^{\lambda(2-\delta)}+(1-\eta)^{r},$
@@
$\leq\mathbb{E}\left[\sum_{\begin{subarray}{c}P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}|\hat{f}(\boldsymbol{k})|^{2}\right]$
$=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]|\hat{f}(\boldsymbol{k})|^{2}$
$\leq c\frac{r}{M}\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}|\hat{f}(\boldsymbol{k})|^{2}\log(1+\|\boldsymbol{k}\|_{\infty})$
$\leq c\frac{r}{M}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}(|\hat{f}(\boldsymbol{k})|r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}\right)$
×$\qquad\times\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\leq c\frac{r}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}^{2}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}.$
@@
$\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$ $\leq\frac{1}{\alpha\delta}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{\|\boldsymbol{k}\|_{\infty}^{\alpha\delta}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\leq\frac{1}{\alpha\delta}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2-\delta}}$
$\leq\frac{1}{\alpha\delta}\left(B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\right)^{2-\delta}.$
@@
$\leq c\frac{r}{\alpha\delta M}\left(B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\right)^{2-\delta}+(1-\eta)^{r}$
$\leq\frac{rC_{\lambda,\delta}}{\alpha(1-\eta)^{\lambda(2-\delta)}M^{\lambda(2-\delta)+1}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}\left(2\zeta(\alpha/\lambda)\right)^{|u|}\right)^{\lambda(2-\delta)}+(1-\eta)^{r},$
@@
Let $M\in\mathbb{N}$, with $M\geq 2$, be given and assume that $M$ and $\eta\in(0,1)$ satisfy (3) for all $1/2\leq\lambda<\alpha$. Let
@@
Then $e^{\mathrm{rms}}(H_{\alpha,\boldsymbol{\gamma}};\tilde{A}^{\mathrm{ran}}_{r,M})$ decays with the order $M^{-\lambda-1/2-\lambda\delta/2}\log M$ for any $1/2\leq\lambda<\alpha$ and $0<\delta\leq 1/\alpha$, which is arbitrarily close to $M^{-\alpha-1/2}$ when $\lambda\to\alpha-0$ and $\delta\to 0+0$. Moreover, $e^{\mathrm{rms}}(H_{\alpha,\boldsymbol{\gamma}};\tilde{A}^{\mathrm{ran}}_{r,M})$ is bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$.
@@ 3.3. Additional remarks
In Algorithm 3.1, we first randomly draw a single $N$, followed by $\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{r}$ conditioned on $N$. Even if this process were replaced by independently and randomly drawing $N_{1},\ldots,N_{r}$ from the set $\mathcal{P}_{M}$ first, and then each $\boldsymbol{z}_{j}$ conditioned on $N_{j}$, a randomized lattice rule based on the pair $(N^{*},\boldsymbol{z}^{*})$ that minimizes the worst-case error $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}},P_{N_{j},\boldsymbol{z}_{j}})$ among $(N_{1},\boldsymbol{z}_{1}),\ldots,(N_{r},\boldsymbol{z}_{r})$ can still yield similar upper bounds on both the (worst-case) randomized error and the RMSE.
@@
As mentioned in Lemma 2.4, computing $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}},P_{N,\boldsymbol{z}_{j}})$ requires a computational cost of $O(dM)$ for each $j\in\{1,\ldots,r\}$. Thus, the necessary cost to run Algorithm 3.1 once is $O(rdM)$. To achieve nearly the optimal rate of the two error criteria with our new randomized rule, it suffices to set
[…]
as discussed in Corollaries 3.5 and 3.7, respectively. This results in a slightly higher cost of $O(\alpha dM\log M)$ compared to the randomized CBC algorithm in [5]. However, by exploiting parallel computation for computing $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}},P_{N,\boldsymbol{z}_{j}})$ for all $j$, the overall cost of running Algorithm 3.1 once remains $O(dM)$, making it slightly more efficient than the randomized CBC algorithm.
@@
[…]
Applying Jensen’s inequality $(\sum_{j}a_{j})^{\lambda}\leq\sum_{j}a_{j}^{\lambda}$, which holds for any $a_{1},a_{2},\ldots>0$ and $10<\lambda<$ 1, to the worst-case error expression in Lemma 2.4, we obtain
$\left(e^{\mathrm{wor}}(H_{\beta,\boldsymbol{\gamma}^{\prime}};Q_{P_{N,\boldsymbol{z}}})\right)^{2\alpha/\beta}$ $\leq\sum_{\boldsymbol{k}\in P_{N,\boldsymbol{z}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{(r_{\beta,\boldsymbol{\gamma}^{\prime}}(\boldsymbol{k}))^{2\alpha/\beta}}$
$=\sum_{\boldsymbol{k}\in P_{N,\boldsymbol{z}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}=\left(e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};Q_{P_{N,\boldsymbol{z}}})\right)^{2}.$
@@
for any $d\in\mathbb{N}$. By choosing $d\propto\lceil M^{\alpha/\beta}\rceil$, the additional term arising from the concatenation of random points does not dominate, ensuring that the rate of convergence remains nearly optimal, regardless of how large $s$ is.

### v3

@@ 3.2. Error bounds
$\leq c\frac{r}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}\left(c^{\prime}_{\lambda,\delta}\left(H_{M,\eta,\alpha,\boldsymbol{\gamma}}\right)^{2-(2\delta+1)/\lambda}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\right)^{1/2}$
$\leq c\frac{r{c^{\prime}_{\lambda,\delta}}^{1/2}}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}\left(\frac{2}{(1-\eta)M}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\right)^{\lambda-(2\delta+1)/2}$
×$\qquad\qquad\times\left(\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\right)^{1/2}$
$=\|f\|_{\alpha,\boldsymbol{\gamma}}\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{\lambda-\delta},$
@@
so that, by applying Cauchy–Schwarz inequality, we obtain
@@
$\leq(1-\eta)^{r}\|f\|_{\alpha,\boldsymbol{\gamma}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{2}\zeta(2\alpha)\right)^{1/2}.$
@@
$e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ $\leq\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{\lambda-\delta}$
$\qquad+(1-\eta)^{r}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{2}\zeta(2\alpha)\right)^{1/2},$
@@
Let $M\in\mathbb{N}$, with $M\geq 2$, be given and assume that $M$ and $\eta\in(0,1)$ satisfy (3). Let
@@
Then $e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ decays with the order $M^{-\lambda+\delta-1/2}\log M$ for any $1/2<\lambda<\alpha$ and $0<\delta<\min(\lambda-1/2,1)$, which is arbitrarily close to $M^{-\alpha-1/2}$ when $\lambda\to\alpha^{-}$ and $\delta\to 0^{+}$. Moreover, $e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ is bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$.
@@
Let $\alpha>1/2$ be a real number and $\boldsymbol{\gamma}\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ be a sequence of non-negative weights, with each $\gamma_{j}$ bounded above by $1$. Furthermore, let $r,M\in\mathbb{N}$, with $M\geq 2$, be given. Assume that $M$ and $\eta\in(0,1)$ satisfy (3). Then it holds that
@@
$\leq\frac{rC_{\lambda,\delta}}{\alpha(1-\eta)^{\lambda(2-\delta)}M^{\lambda(2-\delta)+1}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{\lambda(2-\delta)}+(1-\eta)^{r},$
@@
$\leq\mathbb{E}\left[\sum_{\begin{subarray}{c}\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}|\hat{f}(\boldsymbol{k})|^{2}\right]$
$=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]|\hat{f}(\boldsymbol{k})|^{2}$
$\leq c\frac{r}{M}\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}|\hat{f}(\boldsymbol{k})|^{2}\log(1+\|\boldsymbol{k}\|_{\infty})$
$\leq c\frac{r}{M}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}(|\hat{f}(\boldsymbol{k})|r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}\right)\times\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\leq c\frac{r}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}^{2}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}.$
@@
$\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{\log(1+\|\boldsymbol{k}\|_{\infty})}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$ $\leq\frac{1}{\alpha\delta}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{\|\boldsymbol{k}\|_{\infty}^{\alpha\delta}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\leq\frac{1}{\alpha\delta}\sup_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2-\delta}}$
$\leq\frac{1}{\alpha\delta}\left(H_{M,\eta,\alpha,\boldsymbol{\gamma}}\right)^{2-\delta}.$
@@
$\leq c\frac{r}{\alpha\delta M}\left(H_{M,\eta,\alpha,\boldsymbol{\gamma}}\right)^{2-\delta}+(1-\eta)^{r}$
$\leq\frac{rC_{\lambda,\delta}}{\alpha(1-\eta)^{\lambda(2-\delta)}M^{\lambda(2-\delta)+1}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{\lambda(2-\delta)}+(1-\eta)^{r},$
@@
Let $M\in\mathbb{N}$, with $M\geq 2$, be given and assume that $M$ and $\eta\in(0,1)$ satisfy (3). Let
@@
Then $e^{\mathrm{rms}}(H_{\alpha,\boldsymbol{\gamma}};\tilde{A}^{\mathrm{ran}}_{r,M})$ decays with the order $M^{-\lambda-1/2-\lambda\delta/2}\log M$ for any $1/2\leq\lambda<\alpha$ and $0<\delta\leq 1/\alpha$, which is arbitrarily close to $M^{-\alpha-1/2}$ when $\lambda\to\alpha^{-}$ and $\delta\to 0^{+}$. Moreover, $e^{\mathrm{rms}}(H_{\alpha,\boldsymbol{\gamma}};\tilde{A}^{\mathrm{ran}}_{r,M})$ is bounded independently of the dimension $d$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$.
@@ 3.3. Additional remarks
In Algorithm 3.1, we first randomly draw a single $N$, followed by $\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{r}$ conditioned on $N$. Even if this process were replaced by independently and randomly drawing $N_{1},\ldots,N_{r}$ from the set $\mathcal{P}_{M}$ first, and then each $\boldsymbol{z}_{j}$ conditioned on $N_{j}$, a randomized lattice rule based on the pair $(N^{*},\boldsymbol{z}^{*})$ that minimizes the worst-case error $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};P_{N_{j},\boldsymbol{z}_{j}})$ among $(N_{1},\boldsymbol{z}_{1}),\ldots,(N_{r},\boldsymbol{z}_{r})$ can still yield similar upper bounds on both the (worst-case) randomized error and the RMSE.
@@
As mentioned in Lemma 2.4, computing $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};P_{N,\boldsymbol{z}_{j}})$ requires a computational cost of $O(dM)$ for each $j\in\{1,\ldots,r\}$. Thus, the necessary cost to run Algorithm 3.1 once is $O(rdM)$. To achieve nearly the optimal rate of the two error criteria with our new randomized rule, it suffices to set
$r=\left\lceil-\left(\alpha+\frac{1}{2}\right)\frac{\log M}{\log(1-\eta)}\right\rceil\quad\text{or}\quad r=\left\lceil-\left(2\alpha+1\right)\frac{\log M}{\log(1-\eta)}\right\rceil,$
as discussed in Corollaries 3.5 and 3.7, respectively. This results in a cost of $O(\alpha dM\log M)$, which has an additional factor $\alpha$ compared to the cost of the randomized CBC algorithm in [5]. However, by exploiting parallel computation for computing $e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};P_{N,\boldsymbol{z}_{j}})$ for all $j$, the overall cost of running Algorithm 3.1 once remains $O(dM)$, if $O(\log M)$ computational nodes are available. It should be noted that the randomized CBC algorithm in [5] requires a cost of $O(dM\log M)$, using the fast Fourier transform. The $\log M$ factor in the CBC algorithm is not easily parallelizable, which makes our approach slightly more efficient in a parallel computation setting.
@@
$\beta>\alpha>\frac{1}{2}\quad\text{and}\quad\gamma^{\prime}_{j}=\gamma_{j}^{\beta/\alpha}\quad\text{for all $j$.}$
Since the $\ell_{p}$-norm for an infinite sequence is decreasing in $p\geq 1$, we can infer
$\left(e^{\mathrm{wor}}(H_{\beta,\boldsymbol{\gamma}^{\prime}};Q_{P_{N,\boldsymbol{z}}})\right)^{2}$ $=\sum_{\boldsymbol{k}\in P_{N,\boldsymbol{z}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{(r_{\beta,\boldsymbol{\gamma}^{\prime}}(\boldsymbol{k}))^{2}}=\sum_{\boldsymbol{k}\in P_{N,\boldsymbol{z}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2\beta/\alpha}}$
$\leq\left(\sum_{\boldsymbol{k}\in P_{N,\boldsymbol{z}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}\right)^{\beta/\alpha}=\left(e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};Q_{P_{N,\boldsymbol{z}}})\right)^{2\beta/\alpha}.$
@@
for any $d\in\mathbb{N}$. By choosing $d=c\cdot\lceil M^{\alpha/\beta}\rceil$ for some positive constant $c$, the additional term arising from the concatenation of random points does not dominate, ensuring that the rate of convergence remains nearly optimal, regardless of how large $s$ is.